40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 22 22:50:13 2019, gautam, Update, VAC, Pumpdown prep 
    Reply  Thu May 23 15:30:56 2019, gautam, Update, VAC, Pumpdown underway - so far so good! pumpdownPres.png
       Reply  Fri May 24 11:47:15 2019, gautam, Update, VAC, IFO is almost at nominal vacuum pumpdown_20190524.png
          Reply  Mon May 27 11:37:13 2019, gautam, Update, VAC, c1vac is unresponsive 
             Reply  Tue May 28 09:51:33 2019, gautam, Update, VAC, c1vac hard-rebooted Screen_Shot_2019-05-31_at_12.44.54_PM.png
Message ID: 14631     Entry time: Wed May 22 22:50:13 2019     Reply to this: 14634
Author: gautam 
Type: Update 
Category: VAC 
Subject: Pumpdown prep 

I did the following:

  1. Checked the ETMY OSEM sensing matrix and OSEM actuation matrix - more on this later, but everything seems much more reasonable than it was prior to this vent.
  2. Checked that the IMC could be locked with the low-power beam
  3. Aligned the Y-arm cavity using the green beam. Then tweaked the TT1/TT2 alignment until I saw IR flashes in TRY.
  4. Repeated #2 for the X arm, using the BS to control the beam pointing.
  5. Confirmed that the AS beam makes it out of the vacuum. It is only ~30uW in a large (~1cm dia) beam, so not the clearest spot on an IR card, but looks pretty clean, no evidence of clipping. I removed an ND filter on the AS port camera in order to better see the beam on the CRT monitor, this should be re-installed prior to ramping the input power to the IMC again.
  6. With the PRM aligned, I confirmed that I could see resonant flashes in the POP QPD.
  7. With the SRM aligned, I confirmed that I could see SRC cavity flashes on the AS camera.

I think this completes the pre-pumpdown alignment checks we usually do. The detailed plan for tomorrow is here: please have a look and lmk if I missed something. 

ELOG V3.1.3-